# GraphMoE: Amplifying Cognitive Depth of Mixture-of-Experts Network via Introducing Self-Rethinking Mechanism

### v2

@@ Abstract
Traditional Mixture-of-Experts (MoE) networks benefit from utilizing multiple smaller expert models as opposed to a single large network. However, these experts typically operate independently, leaving a question open about whether interconnecting these models could enhance the performance of MoE networks. In response, we introduce GraphMoE, a novel method aimed at augmenting the cognitive depth of language models via a self-rethinking mechanism constructed on Pseudo Graph MoE networks. GraphMoE employs a recurrent routing strategy to simulate iterative thinking steps, thereby facilitating the flow of information among expert nodes. We implement the GraphMoE architecture using Low-Rank Adaptation techniques (LoRA) and conduct extensive experiments on various benchmark datasets. The experimental results reveal that GraphMoE outperforms other LoRA based models, achieving state-of-the-art (SOTA) performance. Additionally, this study explores a novel recurrent routing strategy that may inspire further advancements in enhancing the reasoning capabilities of language models.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by the GPT series, have dramatically advanced the field of Natural Language Processing (NLP), demonstrating exceptional performance across a variety of tasks including commonsense reasoningYang et al. (2024), creative generationLoakman et al. (2023), dialogue generationLv et al. (2024); Zhao et al. (2023, 2024b), and summarizationGoldsack et al. (2023); Zhao et al. (2024a). In recent years, there has been a growing interest in refining LLM architectures to deliver superior performance with reduced parameters and lower GPU memory consumptionTang et al. (2023). A prominent strategy in this endeavor is the Mixture-of-Experts (MoE) architecture Cai et al. (2024), which pre-trains multiple Feed-Forward Networks (FFNs) as specialized experts, activating a select few during inference. This strategy allows experts to operate optimally within their domains of expertise, providing enhanced performance compared to LLMs with a similar number of activated parameters.
@@
However, expert models are employed independently through a linear routing strategy. Given that these models are optimized for distinct input distributions, we hypothesize that fostering collaboration among expert models akin to connected nodes in a graph network may further exploit their problem-solving capabilities. Inspired by the knowledge aggregation on pseudo graph Tang et al. (2023), we propose employing a recurrent routing strategy wherein expert models function as graph nodes. We posit that with increased graphical message transmission between expert nodes, these models can progressively narrow the semantic gap between the MoE model outputs and human references. Thus, we introduce GraphMoE(Graph Mixture of Experts), a pioneering approach aimed at deepening the cognitive capabilities of language models by integrating a self-rethinking mechanism into constructed pseudo-graph MoE networks. This methodology may bolster the reasoning capabilities of LLMs by enabling incremental problem-solving, akin to the Chain-of-Thoughts (CoT) strategy.
@@
GraphMoE emulates human-like iterative reasoning by allowing the model to continuously revisit and refine its comprehension of input data through multiple reasoning cycles. This is accomplished by implementing a recurrent routing strategy on a pseudo-graph formed by expert models. These models transmit their outputs as signals for selecting subsequent expert batches and as inputs encapsulating aggregated graph features from prior reasoning rounds—a mechanism we term the “self-rethinking mechanism”. Given the challenges inherent in pre-training a MoE LLM from scratch, we chose to implement the GraphMoE architecture utilizing Low-Rank Adaptation (LoRA) techniques [^1] As stated in §2.1, LoRA introduces a set of additional, small-scale training parameters to fine-tune a LLM for a downstream task, while keeping most of the LLM’s original parameters frozen., enabling evaluation of GraphMoE via Supervised Fine-Tuning (SFT) across several benchmarks.
@@ 2.3 Mixture of Experts (MoE)
Numerous studies have investigated the application of Parameter-Efficient Fine-Tuning (PEFT) methods to introduce additional trainable parameters for implementing pseudo MoE structures within LLMs (Dou et al., 2024; Luo et al., 2024; Gao et al., 2024; Li et al., 2024; Gou et al., 2023). These models expand the conventional single feed-forward network (FFN) architecture and its corresponding representation space into multiple subspaces, thereby effectively emulating the behavior of MoE architectures. Prominent examples include MoLA (Gao et al., 2024), LoRAMoE (Dou et al., 2024), and MixLoRA (Li et al., 2024), all of which have demonstrated SOTA performance relative to other PEFT-based methods across various benchmarks. In this study, we augment MoE architectures by enhancing their reasoning depth and demonstrate the efficacy of our novel architecture when integrated with PEFT-based LLM baselines.
Figure 2: Overview of GraphMoE architecture.
@@ 3.1 MoE Transformation
Additionally, we introduce a trainable linear layer $\mathbf{R}^{\ell}\in\mathbb{R}^{n\times d}$, referred to as the Router, where $n$ denotes the total number of experts, and $d$ represents the dimension of the hidden state. The Router determines the importance of each expert during forward propagation by computing a relevance score for every expert. Based on the Router’s output, we select the top $k$ experts and normalize their corresponding weights to ensure effective routing, as described in Equations 3 and 4.
@@
=$=\begin{cases}\mathbf{\hat{s}}^{\ell}\left[i\right],&\text{if }\mathbf{\hat{s}}^{\ell}\left[i\right]\text{ is in the top }k,\\ 0,&\text{otherwise.}\end{cases}$
@@ 3.2 Pseudo Reasoning Graph Construction
5      $\mathbf{\tilde{s}}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{N})$leftarrow $\operatorname{Softmax}\left(\mathbf{R}^{\ell}*\mathbf{x}^{\ell}_{t}\right)$;
@@
7      $\mathbf{s}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{N})$leftarrow $\operatorname{Top-}k\left(\mathbf{\tilde{s}}^{\ell}_{1}\right)$;
@@
9      for $i$ in {selected experts $\left\{\text{selected experts }E\right\}$ do
10             $\tilde{\mathbf{W}}_{\text{G}_{i}}^{\ell}$: $(\mathtt{D^{\prime}},\mathtt{D})$leftarrow $\mathbf{W}_{\text{G}_{i}}^{\ell}+\mathbf{B}_{\text{G}_{i}}^{\ell}*\mathbf{A}_{\text{G}_{i}}^{\ell}$;
@@
14            $\hat{\mathbf{c}}^{\text{G}}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$leftarrow $\tilde{\mathbf{W}}_{\text{G}_{i}}^{\ell}*\mathbf{x}_{t}^{\ell}$;
@@
16            $\hat{\mathbf{c}}^{\text{U}}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$leftarrow $\tilde{\mathbf{W}}_{\text{U}_{i}}^{\ell}*\mathbf{x}_{t}^{\ell}$;
@@
18            $\hat{\mathbf{c}_{i}}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$leftarrow $\sigma\left(\hat{\mathbf{c}}^{\text{G}}_{i}\right)\odot\hat{\mathbf{c}}^{\text{U}}_{i}$;
@@
20            $\mathbf{c}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$leftarrow $\tilde{\mathbf{W}}_{\text{D}_{i}}^{\ell}*\hat{\mathbf{c}_{i}}$;
@@
22            $\mathbf{y}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$leftarrow $\mathbf{y}^{\ell}_{t}+\mathbf{s}^{\ell}_{t}\left[:,i\right]\odot\mathbf{c}_{i}$;
@@
27             $\mathbf{z}_{t}^{\ell}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$leftarrow $\sigma\left(\mathbf{W}^{\ell}_{z}*\left[\mathbf{h}^{\ell}_{t-1},\mathbf{y}^{\ell}_{t}\right]\right)$;
@@
29            $\mathbf{r}_{t}^{\ell}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$leftarrow $\sigma\left(\mathbf{W}^{\ell}_{r}*\left[\mathbf{h}^{\ell}_{t-1},\mathbf{y}^{\ell}_{t}\right]\right)$;
@@
31            $\mathbf{\hat{h}}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}}+\mathtt{D})$leftarrow $\left[\mathbf{r}_{t}^{\ell}\odot\mathbf{h}_{t-1}^{\ell},\mathbf{\mathbf{x}}_{t}^{\ell}\right]$;
@@
33            $\mathbf{h}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$leftarrow $\sigma\left(\mathbf{W}^{\ell}_{o}*\mathbf{\hat{h}}^{\ell}_{t}+\mathbf{b}_{o}^{\ell}\right)$;
@@
35            $\mathbf{h}^{\ell}_{t}$leftarrow $\left(\mathbf{1}-\mathbf{z}_{t}^{\ell}\right)\odot\mathbf{h}^{\ell}_{t-1}+\mathbf{z}_{t}^{\ell}\odot\mathbf{h}_{t}^{\ell}$;
@@
37            $\mathbf{g}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$leftarrow $\mathbf{W}_{g}^{\ell}\cdot\mathbf{h}^{\ell}_{t}$;
@@
40            $\mathbf{x}^{\ell}_{t+1}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$leftarrow $\mathbf{x}^{\ell}_{t}+\mathbf{g}^{\ell}_{t}$
@@ 3.3 GraphMoE Training
where $\text{where }f_{i}$ $=\frac{1}{|\mathcal{M}|}\sum_{\left(\mathbf{x},\mathbf{y}\right)\in\mathcal{M}}\mathds{1}\{\mathbf{s}_{\mathbf{x}}\left[i\right]>0\}\text{,}$
@@ 4.4 Implementation Details
In the following experiments utilizing the GraphMoE framework, three additional hyperparameters are configured as follows: (1) The number of iterations for applying the GRU gate within the recurrent routing, referred to as the reasoning (self-rethinking) round $T$; our GraphMoE models are configured with a default setting of $3$. During the initial round, the outputs of expert models are obtained without the GRU gate’s feature transmission. (2) The GRU network is set to a single layer. (3) The GRU hidden size scale is predefined as $0.1\times 4096$, where $4096$ represents the model dimension of the LLaMA-3-8B model. (4) TThe hyperparameter $\lambda$ associated with the auxiliary load balance training loss $\mathcal{L}_{LB}$ is configured to be $0.01$.

### v3

@@ Abstract
Traditional Mixture-of-Experts (MoE) networks benefit from utilizing multiple smaller expert models as opposed to a single large network. However, these experts typically operate independently, leaving a question open about whether interconnecting these models could enhance the performance of MoE networks. In response, we introduce GraphMoE, a novel method aimed at augmenting the cognitive depth of language models via a self-rethinking mechanism constructed on Pseudo Graph MoE networks. GraphMoE employs a recurrent routing strategy to simulate iterative thinking steps, thereby facilitating the flow of information among expert nodes. We implement the GraphMoE architecture using Low-Rank Adaptation techniques (LoRA) and conduct extensive experiments on various benchmark datasets. The experimental results reveal that GraphMoE outperforms other LoRA based models, achieving state-of-the-art (SOTA) performance. Additionally, this study explores a novel recurrent routing strategy that may inspire further advancements in enhancing the reasoning capabilities of language models. Our code is available at https://github.com/fan2goa1/GraphMoE_raw.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by the GPT series, have dramatically advanced the field of Natural Language Processing (NLP), demonstrating exceptional performance across a variety of tasks including commonsense reasoning Yang et al. (2024), creative generation Loakman et al. (2023), dialogue generation Lv et al. (2024); Zhao et al. (2023, 2024b), and summarization Goldsack et al. (2023); Zhao et al. (2024a). In recent years, there has been a growing interest in refining LLM architectures to deliver superior performance with reduced parameters and lower GPU memory consumption Tang et al. (2023a). A prominent strategy in this endeavor is the Mixture-of-Experts (MoE) architecture Cai et al. (2024), which pre-trains multiple Feed-Forward Networks (FFNs) as specialized experts, activating a select few during inference. This strategy allows experts to operate optimally within their domains of expertise, providing enhanced performance compared to LLMs with a similar number of activated parameters.
@@
However, expert models are employed independently through a linear routing strategy. Given that these models are optimized for distinct input distributions, we hypothesize that fostering collaboration among expert models akin to connected nodes in a graph network may further exploit their problem-solving capabilities. Inspired by the knowledge aggregation on pseudo graph Tang et al. (2023a), we propose employing a recurrent routing strategy wherein expert models function as graph nodes. We posit that with increased graphical message transmission between expert nodes, these models can progressively narrow the semantic gap between the MoE model outputs and human references. Thus, we introduce GraphMoE (Graph Mixture of Experts), a pioneering approach aimed at deepening the cognitive capabilities of language models by integrating a self-rethinking mechanism into constructed pseudo-graph MoE networks. This methodology may bolster the reasoning capabilities of LLMs by enabling incremental problem-solving, akin to the Chain-of-Thoughts (CoT) strategy.
@@
GraphMoE emulates human-like iterative reasoning by allowing the model to continuously revisit and refine its comprehension of input data through multiple reasoning cycles [^1] This motivation is based on the dual process theory in cognitive science, constructing a pseudo-graph iteratively by integrating a base model for direct inference (System 1) with a LoRA-based rethinking module for deliberate reasoning (System 2). The graph architecture is particularly apt for creating such an iterative thinking system, and related implementations can be found in Ding et al. (2019); Yan et al. (2023).. This is accomplished by implementing a recurrent routing strategy on a pseudo-graph formed by expert models. These models transmit their outputs as signals for selecting subsequent expert batches and as inputs encapsulating aggregated graph features from prior reasoning rounds—a mechanism we term the “self-rethinking mechanism”. Given the challenges inherent in pre-training a MoE LLM from scratch, we chose to implement the GraphMoE architecture utilizing Low-Rank Adaptation (LoRA) techniques [^2] As stated in §2.1, LoRA introduces a set of additional, small-scale training parameters to fine-tune a LLM for a downstream task, while keeping most of the LLM’s original parameters frozen., enabling evaluation of GraphMoE via Supervised Fine-Tuning (SFT) across several benchmarks.
@@ 2.3 Mixture of Experts (MoE)
Numerous studies have investigated the application of Parameter-Efficient Fine-Tuning (PEFT) methods to introduce additional trainable parameters for implementing pseudo MoE structures within LLMs (Dou et al., 2024; Luo et al., 2024; Gao et al., 2024; Li et al., 2024; Gou et al., 2023). These models expand the conventional single feed-forward network (FFN) architecture and its corresponding representation space into multiple subspaces, thereby effectively emulating the behavior of MoE architectures.
Prominent examples include MoLA (Gao et al., 2024), LoRAMoE (Dou et al., 2024), and MixLoRA (Li et al., 2024), all of which have demonstrated state-of-the-art performance compared to other PEFT-based methods across various benchmarks. Consequently, we integrate these three distinct LoRA MoE methodologies into our GraphMoE framework. The primary differences between these models are as follows: LoRAMoE incorporates vanilla attention LoRA layers and MLP plugged with LoRA-MoE layers; MoLA integrates plugged LoRA-MoE layers in both attention and MLP layers; and MixLoRA combines fused LoRA-MoE modules in attention and MLP layers. Further analysis of these LoRA-MoE implementations can be found in Li et al. (2024).
In this study, we augment MoE architectures by enhancing their reasoning depth and demonstrate the efficacy of our novel architecture when integrated with PEFT-based LLM baselines.
Figure 2: Overview of GraphMoE architecture. In this figure, the original Feed-Forward Network (FFN) layer in each transformer block is modified. FFN is also known as a Multi-Layer Perceptron (MLP).
@@ 3 Method
The overall framework is illustrated in Figure 2. The self-rethinking mechanism operates within the GraphMoE Feed-Forward Network (FFN) layer, which replaces the original FFN layer found in standard transformer-based LLMs. In this diagram, the LoRA MoE is derived from the MixLoRA architecture. When using alternative architectures such as MoLA and LoRAMoE, the illustration of the Expert component varies. Nonetheless, the implementation of the self-rethinking module remains consistent, as it incorporates an additional GRU model introduced in §3.2.
@@ 3.1 MoE Transformation
Additionally, we introduce a trainable linear layer $\mathbf{R}^{\ell}\in\mathbb{R}^{n\times d}$, referred to as the Router, where $n$ denotes the total number of experts, and $d$ represents the dimension of the hidden state. The Router determines the importance of each expert during forward propagation by computing a relevance score for every expert. Based on the Router’s output, we select the top $k$ experts and normalize their corresponding weights to ensure effective routing, as described in Equations 3 and 4. In subsequent experiments, unless stated otherwise, our setup activates 2 out of 8 experts, thus setting $k$ to 2.
@@
[…]
@@ 3.2 Pseudo Reasoning Graph Construction
5 $\mathbf{\tilde{s}}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{N})$ $\leftarrow$ $\operatorname{Softmax}\left(\mathbf{R}^{\ell}*\mathbf{x}^{\ell}_{t}\right)$;
@@
7 $\mathbf{s}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{N})$ $\leftarrow$ $\operatorname{Top-}k\left(\mathbf{\tilde{s}}^{\ell}_{1}\right)$;
@@
9 for $i$ in $\left\{\text{selected experts }E\right\}$ do
10    $\tilde{\mathbf{W}}_{\text{G}_{i}}^{\ell}$: $(\mathtt{D^{\prime}},\mathtt{D})$ $\leftarrow$ $\mathbf{W}_{\text{G}_{i}}^{\ell}+\mathbf{B}_{\text{G}_{i}}^{\ell}*\mathbf{A}_{\text{G}_{i}}^{\ell}$;
@@
14    $\hat{\mathbf{c}}^{\text{G}}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$ $\leftarrow$ $\tilde{\mathbf{W}}_{\text{G}_{i}}^{\ell}*\mathbf{x}_{t}^{\ell}$;
@@
16    $\hat{\mathbf{c}}^{\text{U}}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$ $\leftarrow$ $\tilde{\mathbf{W}}_{\text{U}_{i}}^{\ell}*\mathbf{x}_{t}^{\ell}$;
@@
18    $\hat{\mathbf{c}_{i}}$: $(\mathtt{B},\mathtt{P},\mathtt{D^{\prime}})$ $\leftarrow$ $\sigma\left(\hat{\mathbf{c}}^{\text{G}}_{i}\right)\odot\hat{\mathbf{c}}^{\text{U}}_{i}$;
@@
20    $\mathbf{c}_{i}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$ $\leftarrow$ $\tilde{\mathbf{W}}_{\text{D}_{i}}^{\ell}*\hat{\mathbf{c}_{i}}$;
@@
22    $\mathbf{y}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$ $\leftarrow$ $\mathbf{y}^{\ell}_{t}+\mathbf{s}^{\ell}_{t}\left[:,i\right]\odot\mathbf{c}_{i}$;
@@
27    $\mathbf{z}_{t}^{\ell}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$ $\leftarrow$ $\sigma\left(\mathbf{W}^{\ell}_{z}*\left[\mathbf{h}^{\ell}_{t-1},\mathbf{y}^{\ell}_{t}\right]\right)$;
@@
29    $\mathbf{r}_{t}^{\ell}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$ $\leftarrow$ $\sigma\left(\mathbf{W}^{\ell}_{r}*\left[\mathbf{h}^{\ell}_{t-1},\mathbf{y}^{\ell}_{t}\right]\right)$;
@@
31    $\mathbf{\hat{h}}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}}+\mathtt{D})$ $\leftarrow$ $\left[\mathbf{r}_{t}^{\ell}\odot\mathbf{h}_{t-1}^{\ell},\mathbf{\mathbf{x}}_{t}^{\ell}\right]$;
@@
33    $\mathbf{h}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\overline{\mathtt{D}})$ $\leftarrow$ $\sigma\left(\mathbf{W}^{\ell}_{o}*\mathbf{\hat{h}}^{\ell}_{t}+\mathbf{b}_{o}^{\ell}\right)$;
@@
35    $\mathbf{h}^{\ell}_{t}$ $\leftarrow$ $\left(\mathbf{1}-\mathbf{z}_{t}^{\ell}\right)\odot\mathbf{h}^{\ell}_{t-1}+\mathbf{z}_{t}^{\ell}\odot\mathbf{h}_{t}^{\ell}$;
@@
37    $\mathbf{g}^{\ell}_{t}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$ $\leftarrow$ $\mathbf{W}_{g}^{\ell}\cdot\mathbf{h}^{\ell}_{t}$;
@@
40    $\mathbf{x}^{\ell}_{t+1}$: $(\mathtt{B},\mathtt{P},\mathtt{D})$ $\leftarrow$ $\mathbf{x}^{\ell}_{t}+\mathbf{g}^{\ell}_{t}$
@@
To enhance clarity in illustrating the long processing pipeline, we summarize the entire methodology in Algorithm 1. The process involves $T$ reasoning iterations, referred to as self-rethinking. Initially, features from the FFN layer of transformer blocks are derived via a LoRA MoE forward pass. Subsequently, these features are transmitted to the subsequent Low-Rank GRU module, which computes both the residual connection within the pseudo graph and the merged feature for the input of the succeeding iteration.
@@ 3.3 GraphMoE Training
$\text{where }f_{i}$ $=\frac{1}{|\mathcal{M}|}\sum_{\left(\mathbf{x},\mathbf{y}\right)\in\mathcal{M}}\mathds{1}\{\mathbf{s}_{\mathbf{x}}\left[i\right]>0\}\text{,}$
@@ 4.4 Implementation Details
In the following experiments utilizing the GraphMoE framework, a range of main hyperparameters are configured as follows: (1) The number of iterations for applying the GRU gate within the recurrent routing, referred to as the reasoning (self-rethinking) round $T$; our GraphMoE models are configured with a default setting of $3$. During the initial round, the outputs of expert models are obtained without the GRU gate’s feature transmission. (2) The GRU network is set to a single layer. (3) The GRU hidden size scale is predefined as $0.1\times 4096$ [^5] The actual hidden size is approximated to the nearest integer using Python fucntion “int()”., where $4096$ represents the model dimension of the LLaMA-3-8B model. (4) The hyperparameter $\lambda$ associated with the auxiliary load balance training loss $\mathcal{L}_{LB}$ is configured to be $0.01$. (5) Unless otherwise specified, the expert activation in the MoE architecture is set to a total of 8 experts per MoE module, with 2 experts activated each time. [^6] Additional discussions regarding the implementation details of this paper can be found in Appendix A.